OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 5 2x - 4y = 10 A. Infinitely many solutions B. No solutions C. (0, 0) D. (5, 10)

OpenStudy (gamer56):

Here is what I think: 2*(x-2y)=2*6 2x-4y=12 now 2x-4y=10 There is no solution...

OpenStudy (gamer56):

If I helped you fan and medal please?

OpenStudy (anonymous):

Don't know if I gave you a medal ?

OpenStudy (anonymous):

and thanks

OpenStudy (gamer56):

Fan and thanks